Fermi National Laboratory
 

Accelerator Update for August 4, 2006 - August 9, 2006

 

Monday August 7

The day shift began with the Recycler Ring (RR) stash at 2.9 E10, Pbar stack at 88.5 E10. MiniBooNE and NuMI are taking beam.

Operations established store 4893 at 8:44 AM with an initial luminosity of 28.65 E30.

The low luminosity was due to an A49H separator spark just before scraping.

Operations put the Recycler into shot setup at 8:49 AM to stash antiprotons.

NuMI turned off at 9:19 AM for access to repair a water leak.

Pbar returns to stacking at 9:57 AM.

MiniBooNe turned off at 1:30 PM to investigate problems with their dewatering system.

Operations put the Recycler into shot setup at 2:47 PM to stash antiprotons.

Pbar returns to stacking at 3:06 PM.

NuMI turns back on at 3:50 PM.

Store #4893 was terminated at 4:03 PM.

At 4:24 PM CDF and D0 went in their halls on controlled access.

At 4:46 PM the Pbar target stopped rotating, Pbar stacking is off until experts repair the motor.

MiniBooNe turned back on at 6:22 PM. They found a bad pump and will turn off tomorrow to replace the pump.

Operations put the Recycler into shot setup at 6:59 PM to stash antiprotons.

Transfers were completed at 7:24 PM. Pbar stacking is still off due to problems with the rotating target.

Shot setup begins at 10:41 PM.

Pbar stacking resumed at 11:51 PM.

Tuesday August 8

The midnight shift began with Operations in shot setup, Pbar is stacking, NuMI and MiniBooNE are taking beam.

Operations established store 4895 at 12:44 AM with an initial luminosity of 89.12 E30.

Operations put the Recycler into shot setup at 3:40 AM to stash antiprotons.

Pbar returns to stacking at 3:53 AM.

MiniBooNe turned off at 6:00 AM to replace a dewatering pump.

Operations put the Recycler into shot setup at 11:41 AM to stash antiprotons.

Pbar returns to stacking at 12:20 PM.

MiniBooNe turned back on at 3:35 PM. A pump was replaced on the dewatering system.

Operations put the Recycler into shot setup at 5:31 PM to stash antiprotons.

Pbar returns to stacking at 5:51 PM.

Store #4895 was terminated at 10:01 PM.

Operations put the Recycler into shot setup at 10:27 PM to stash antiprotons.

Pbar returns to stacking at 10:39 PM.

Wednesday August 9

The midnight shift began with the Recycler Ring (RR) stash at 243.4 E10, Pbar stack at 32.8 E10, and with MiniBooNE and NuMI taking beam.

Shot setup begins at 12:37 AM.

Operations established store 4897 at 3:55 AM with an initial luminosity of 144.48 E30. Shot setup was delayed due to high temperature indications on I:LAM42.

The Plan for Wednesday

Keep the store, stack, stash, run beam to NuMI and MiniBooNE.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 08/09/2006    email Fermilab